# Primus: A Pioneering Collection of Open-Source Datasets for Cybersecurity LLM Training

### v2

@@ Appendix B Primus-Nemotron-CC
Table 15: Token counts before and after deduplication for Primus-Nemotron-CC samples (length > 500) across different score bins.
@@
Due to computational constraints, we were unable to include all samples that met the above criteria. Instead, we computed the total number of tokens (for samples with length > 500) within different score ranges, as shown in Tab.15. Given our computing budget, we aimed to limit the 70B model’s pretraining dataset to approximately 10B tokens. As a result, we selected the 0.98 ~ 0.175 score range, which contains 7.6B tokens, for inclusion in Primus-Pretraining. This dataset will also be released.
Appendix C CTI-Bench
@@
Appendix D Training Hyperparameters
@@ D.1 Pre-Training
D.2 Instruction Fine-Tuning
@@
D.3 Reasoning Fine-Tuning
@@
Appendix E Prompts
@@
Appendix F Safety & Toxicity

### v3

@@ Appendix C Primus-Nemotron-CC
Table 18: Token counts before and after deduplication for Primus-Nemotron-$C$ samples (length > 500) across different score bins.
@@
Due to computational constraints, we were unable to include all samples that met the above criteria. Instead, we computed the total number of tokens (for samples with length > 500) within different score ranges, as shown in Tab.18. Given our computing budget, we aimed to limit the 70B model’s pretraining dataset to approximately 10B tokens. As a result, we selected the 0.98 ~ 0.0175 score range, which contains 7.6B tokens, for inclusion in Primus-Pretraining. This dataset will also be released.
Appendix D CTI-Bench
@@
Appendix E Training Hyperparameters
@@ E.1 Pre-Training
E.2 Instruction Fine-Tuning
@@
E.3 Reasoning Fine-Tuning
@@
Appendix F Sample Outputs
In this section, we present an example of the output of Llama-Primus-Merged for a question selected from the CTI-MCQ dataset. We removed the answer choices and prompted the model to respond freely. Then, we compared its response with that of Llama-3.1-8B-Instruct.
Question.
What mitigation technique can help prevent the installation of unauthorized browser extensions as per the MITRE ATT&CK framework?
Llama-Primus-Merged.
As per the MITRE ATT&CK framework, the technique to help prevent the installation of unauthorized browser extensions is T1176: Software Extensions. This technique involves monitoring and restricting the installation of browser extensions to prevent malicious extensions from being installed.
  • Whitelisting: Maintain a list of approved browser extensions (…)
  • Blacklisting: Maintain a list of known malicious browser extensions (…)
  • Group Policy: Use group policy settings to restrict the installation of browser extensions.
  • Browser Settings: Configure browser settings to disable the installation of extensions or to only allow extensions from trusted sources.
  • Regular Security Audits: Regularly audit browser extensions to identify and remove any unauthorized extensions.
By implementing these controls, you can (…)
Llama-3.1-8b-Instruct.
The MITRE ATT&CK framework recommends the following mitigation technique to prevent the installation of unauthorized browser extensions:
  • Restrict Browser Extensions: (…)
  • Use a Secure Browser: (…)
  • Monitor Browser Extensions: (…)
  • Use a Browser Extension Manager: (…)
  • Implement a Least Privilege Policy: (…)
  • Use a Web Application Firewall (WAF): (…)
  • Implement a Secure Configuration: Implement a secure configuration for browser extensions, such as disabling JavaScript (…)
  • Use a Browser Extension Scanner: (…)
  • Implement a Regular Update Policy: (…)
  • Use a Secure Browser Extension Repository: (…)
Note: We’ve only retained key information; "(…)" indicates omitted details.
You can see that Llama-Primus-Merged immediately and correctly references the MITRE ID T1176, and every mitigation it lists maps exactly to the official framework:
  • Whitelisting/Blacklisting aligns with Execution Prevention (M1038)
  • Policy-based restriction of installations implements Limit Software Installation (M1033)
  • Regular audits satisfy the Audit (M1047) mitigation
In contrast, Llama-3.1-8b-Instruct offers a broader set of controls, such as web application firewalls (WAFs), JavaScript disabling, and extension scanners. While these controls are not incorrect, they are not the official ATT&CK mitigations, indicating weaker factual recall of the cybersecurity framework.
Appendix G Prompts
@@
Appendix H Safety & Toxicity